A: If 550 confidence intervals, each having a level of confidence of 96%, were computed for a population mean μ, approximately how many of the intervals would be expected to contain μ?  (Round your answer to the nearest integer.)

B: If 310 confidence intervals, each having a 95% level of confidence, were computed for a population mean μ, about how many of the intervals can be expected to contain μ?  (Your answer will be an integer, of course.)
Answer #1

A) Expected number of confidence interval which contains μ = 550*0.96 = 528

B) Expected number of confidence interval which contains μ = 310*0.95 = 294.5 = 295